OpenStudy (anonymous):

please i need steps for how to solve lx-4l-2=-1

OpenStudy (cwrw238):

first add 2 to both sides of the equation lx-4l = 1 do you know what the vertical lines either side of the x - 4 signify the absolute value ?

OpenStudy (anonymous):

then dont i add 4 to both sides

OpenStudy (cwrw238):

no - next you have to split it up into 2 equations - because the absolute value is the positive value whether x -4 is positive or negative

OpenStudy (anonymous):

x-4=1 and x-4=-1

OpenStudy (cwrw238):

so we get x - 4 = 1 and also x - 4 = -1 solve these 2 equations for x and you have your solutions.

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

then add 4 to the first one to get x=3 and to the second equation to get x=5

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

2 solutions 3 and 5

OpenStudy (cwrw238):

i've got to go . Glad to be of help.

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw